# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ I Introduction
Quantum criticality is in general a property of equilibrium states such as the Gibbs thermal state (which becomes the ground state at zero temperature limit) Sachdev and Safari (2011); Vojta (2003). Preparing such equilibrium states can be challenging in practice, demanding ultra-low temperatures or extremely long preparation times. In addition, criticality-enhanced quantum precision is only achievable around the phase transition point. This requires fine tuning and thus is mostly beneficial in the context of local sensing, where significant prior information about the parameter of interest is available. Beyond criticalities, the sensing capability of many-body systems have also been investigated in their strongly correlated phases, even in the presence of thermal fluctuations Mihailescu et al. (2023, 2024b, 2024a). Moreover, as an alternative approach, one can also use non-equilibrium dynamics of many-body systems for sensing unknown parameters. These setups are more experimental friendly as their initial state can be a simple product state and entanglement is generated naturally during the dynamics. Non-equilibrium dynamics in many-body systems can be induced through different methods, including quantum quench Mitra (2018); Daley et al. (2012); Calabrese and Cardy (2006); Calabrese et al. (2011); Shi et al. (2024); Gietka et al. (2022), measurement quench Pouyandeh et al. (2014); Bayat et al. (2018); Skinner et al. (2019a); Li et al. (2018a), external driving Chu and Telnov (2004); Gritsev and Polkovnikov (2017); Oka and Kitamura (2019) and adiabatic evolution Mihailescu et al. (2024b). Apart from the probe size, in non-equilibrium quantum sensors, time is also a key resource which plays an important role and has to be included in the definition of standard limit and quantum enhanced sensitivity. In general, Fisher information scales as $I\sim t^{\alpha}L^{\beta}$, where $t$ is time, $L$ is the system size and $\alpha$ and $\beta$ are two exponents Ilias et al. (2022). In non-equilibrium sensors, the standard limit is defined as $\alpha{=}\beta{=}1$. Any situation with $\alpha>1$ or $\beta>1$ is called quantum enhanced sensitivity with the special case of Heisenberg limit, defined as $\alpha{=}\beta{=}2$. In certain systems in which the unitary dynamics preserves the number of excitations, the Fisher information may scale with the number of excitations too. In fact, in Ref. Manshouri et al. (2024) it was shown that in excitation-preserving dynamics Fisher information scales as $I\sim t^{\alpha}L^{\beta}N^{\gamma}$, where $N$ is the number of excitations and $\gamma$ is its corresponding exponent. This is an interesting observation which shows how the sensitivity increases from single excitation subspace, i.e. $N{=}1$, to half filling, i.e. $N{=}L/2$. In addition, Ref. Gietka and Ritsch (2023) proposes using spin-orbit coupled quantum gases to overcome the Heisenberg limit and Ref. Gietka et al. (2022) demonstrates that Fisher information can exhibit exponential scaling as a physical effect by quenching dynamics far beyond the critical point. Unlike criticality-based quantum sensors, the criteria for achieving quantum enhanced precision is not yet well-characterized in non-equilibrium quantum many-body probes. For instance, we know that localized phases are not good for sensing He et al. (2023); Yousefjani et al. (2023a, 2024a); Sahoo et al. (2024) but what features in extended phases are crucial for quantum enhanced sensitivity are not yet identified.
@@
In practice, no quantum device is perfectly isolated from its environment and thus decoherence is an inevitable part of any quantum protocol. Consequently, any quantum dynamics is actually an open system evolution. Different formalisms have been developed for addressing open system dynamics, among them are: (i) master equation Carmichael (1993); and (ii) non-Hermitian Hamiltonians Rotter (2009). In the former, the evolution of the system and its environment is described by a unitary operation. By tracing out the environment, the evolution of the system, typically within Born-Markov approximations, is thus characterized by the Lindblad (or Gorini-Kossakowski-Sudarshan-Lindblad) master equation Breuer et al. (2002). In this formalism a quantum system initially prepared in a pure state may indeed evolve to a mixed state due to the action of jump operators — a signature of information leakage (irreversible process) to the environment. In non-Hermitian formalism, the environment is monitored for tracking a trajectory in which no jump operator action takes place Plenio and Knight (1998); Daley (2014). As a result, an initially pure quantum state remains pure throughout the dynamics, despite being evolved effectively as an open quantum system. Other mechanisms that address or profit from dissipation include decoherence-free subspaces Lidar et al. (1998); Duan and Guo (1997), by coupling the quantum resources to a common environment that can be measured at least in part Braun and Martin (2011) and coherent averaging (Fraïsse and Braun, 2015) more generally leading to quantum-enhanced measurements without entanglement (Braun et al., 2018), dynamical decoupling Viola et al. (1999); Khodjasteh and Lidar (2005), and quantum reservoir engineering Poyatos et al. (1996). The latter technique involves adding an engineered Lindbladian term to the open system dynamics. This approach can be designed to steer the system into a desired pure steady-state, effectively cooling the system to that state Myatt et al. (2000); Pielawa et al. (2007, 2010); Groszkowski et al. (2022). Both master equation and non-Hermitian formalisms have been investigated for quantum sensing. Indeed, a general joint (system plus environment) quantum Fisher information upper bound has been formulated Escher et al. (2011a). Hence, one would expect that such open quantum system dynamics may also achieve quantum enhanced sensitivity provided that the system goes through a phase transition, which is again accompanied by an eigenvalue gap closing. In the master equation formalism, the phase transition is determined for the steady state of the system at the point that the eigenvalues of the Liouvillian operator shows gap closing Minganti et al. (2018). An example of such systems with quantum enhanced sensitivity is boundary time crystals Montenegro et al. (2023). In the non-Hermitian systems, quantum enhanced sensitivity has been shown to emerge at the exceptional points Wiersig (2014); Kononchuk et al. (2022), where two or more eigenvalues and corresponding eigenstates coalesce, as well as in the cases of different types of gap closing in the complex energy plane Sarkar et al. (2024).
@@
In this review, we explore the capacity of many-body systems for serving as quantum sensors in both equilibrium and non-equilibrium regimes. Our paper is complementary to a few previous review papers on different aspects of quantum sensing. In Ref. Paris (2009), the theoretical aspects of quantum estimation theory are comprehensively discussed. On the other hand, Ref. Degen et al. (2017) mostly focuses on experimental implementation of quantum sensors. More recent reviews have been dedicated for exploring Fisher information and its properties Meyer (2021); Liu et al. (2020). In addition, recently an essay has also been dedicated to atomic, molecular, and optical platforms for realizing quantum sensors Ye and Zoller (2024). While quantum many-body sensors have received significant attention in recent years, these aforementioned recent progresses have not been covered in the previous review articles. This makes it timely to provide a review on the progress towards quantum many-body sensors.
@@ II Elements of Parameter Estimation
Precise measurements of relevant quantities are essential in science. In physics, high-precision measurements lead us to test the theory of special relativity Einstein (1923, 1905); Michelson and Morley (1887); Rossi and Hall (1941); Hafele and Keating (1972a, b), scrutinize the theory of general relativity via gravitational wave detection Einstein (1923); Phy (2016), and validate the Standard Model of physics by confirming the existence of the elusive Higgs particle Englert and Brout (1964); Higgs (1964); Guralnik et al. (1964); The CMS Collaboration (2022). Thus, pursuing novel schemes and strategies for achieving high-precision measurements are highly desirable for advancing science. However, measuring a relevant quantity is only sometimes a straightforward procedure. Furthermore, quantities such as coupling strengths, temperature, and in general, those lacking corresponding quantum observables, such as entanglement or purity, can only be determined indirectly. Therefore, one requires to infer a relevant quantity of the system by indirect means. The art of inferring such a relevant quantity by means of measuring another is known as parameter estimation Sidhu and Kok (2020); D’ariano et al. (2000); Paris (2009).
@@
Figure 2: Sketch of parameter estimation steps: Standard scheme for estimating an unknown parameter $\theta$. An initial quantum state $\rho$ dynamically encodes the parameter $\theta$ to be estimated as $\rho_{\theta}$. To extract information about $\theta$, a set of measurements via positive operator-valued measure (POVM) $\{\Pi_{x}\}$, are performed on the probe with random outcomes $x$. The resulting measurement outcomes are fed into a classical estimator $\hat{\Theta}$ to infer the unknown parameter $\theta$.
All the above steps need optimization to successfully estimate the parameters. It is important to note that the actual values of the parameters are always unknown by definition, and one must infer them indirectly. One common methodology to quantify the uncertainty in estimating a single unknown parameter $\theta$ is accessing the system of interest through measurements with observable $O$, which leads to estimating the error of $\theta$ using the error propagation formula Sidhu and Kok (2020):
$\delta\theta=\frac{\Delta O}{\left|\partial_{\theta}\langle O\rangle\right|},\hskip 28.45274pt\frac{\partial}{\partial\theta}:=\partial_{\theta}$ (1)
where $\Delta O$ and $\langle O\rangle$ are the standard deviation and the expectation value of the outcomes associated with $O$, respectively. The denominator in Eq. (1) indicates that in order to decrease the error in the estimation of $\theta$ (i.e., $\delta\theta\ll 1$), one needs to increase the denominator, i.e., $\left|\partial_{\theta}\langle O\rangle\right|\gg 1$. To achieve this, it is necessary to find a set of outcomes with an expectation value $\langle O\rangle$ that varies significantly with respect to $\theta$. Conversely, a set of outcomes that varies weakly with $\theta$ (i.e., flat slope $|\partial_{\theta}\langle O\rangle|\ll 1$) will result in a substantial error (i.e., $\delta\theta\gg 1$). While Eq. (1) offers a heuristic explanation of a necessary ingredient for estimating a parameter, a more rigorous and operational framework must be provided to determine a formal figure of merit for quantifying estimation performance. In fact, Eq. (1) does not take into account the entire probability distribution of the outcomes and cannot provide the framework to determine a proper figure of merit for quantifying the performance of an estimation strategy.
@@ II.1 Global and local estimation theory
A suitable strategy to estimate a parameter usually depends on the amount of a priori information one has on the parameter itself. If no a priori information is avalaible, the estimation strategy, i.e. the choice of the measure to be performed on the probe and the procedure employed to process data, should be optimal in average i.e. should be equally good for any value of the parameter. The set of techniques and methods used to choose and optimize the figure of merit in this case is known as global estimation theory Helstrom (1976); Holevo (2011a). Once a global estimation has been performed an experimenter is left with some a priori information to be improved, i.e. one should look for figures of merits that select the best estimation strategy for a specific value or range of values of a parameter. This optimization problem is addressed by local estimation theory Bengtsson and Życzkowski (2017); Van Trees (2004); Helstrom (1976) which allows one to achieve the ultimate bound on precision. The quantum version, namely the local quantum estimation theory Braunstein and Caves (1994); Paris (2009) is the fundamental tool of quantum probing, sensing and metrology and is the main subject of the present review, with the exception of Section XIV, where various aspects of global sensing are discussed.
@@ II.3 Optimizing the classical bound: the QFI
$p(x|\theta)=\text{Tr}[\Pi_{x}\rho_{\theta}].$ (4)
Assuming regularity conditions, and since the trace and the derivative are linear operators, the derivative of the probability distribution $\partial_{\theta}p(x|\theta)$ translates into taking the derivative of the quantum state $\partial_{\theta}\rho_{\theta}$, specifically $\partial_{\theta}p(x|\theta)=\text{Tr}[\Pi_{x}\partial_{\theta}\rho_{\theta}]$. Here, the derivative of the quantum state satisfies the Lyapunov equation given by:
$\partial_{\theta}\rho_{\theta}=\frac{1}{2}\{L_{\theta},\rho_{\theta}\},$ (5)
@@
$I(\theta)\leq I_{Q}(\theta):=\text{Tr}[\rho_{\theta}L_{\theta}^{2}],$ (6)
by incorporating the Schwartz inequality. Here, $I_{Q}(\theta)$ denotes the quantum Fisher information (QFI), leading to the quantum Cramér-Rao bound Paris (2009); Yuen and Lax (1973); Helstrom and Kennedy (1974); Braunstein and Caves (1994); Sidhu and Kok (2020); Albarelli et al. (2020); Liu et al. (2019):
@@
Note that the derivative of the quantum state satisfies other expressions, e.g., the Right Logarithmic Derivative (RLD) Albarelli et al. (2020). This gives rise to different quantum lower bounds. Hence, as opposed to the CFI, the QFI is not unique (see Ref. Meyer (2021) for a thorough discussion and Ref. Albarelli et al. (2020) for different achievable bounds).
The QFI of Eq. (6) accounts for the ultimate precision limit for the estimation of $\theta$. Note that the QFI, as the maximization of the CFI over all possible measurements, depends on the quantum statistical model $\rho_{\theta}$ and the SLD $L_{\theta}$ only, and it does not depend explicitly on a particular measurement basis. The optimal POVM can be composed with the set of projectors over the eigenstates of $L_{\theta}$ with an associated optimal quantum estimatorParis (2009).
To saturate the quantum Cramér-Rao bound in Eq. (2), one needs both the optimal POVM (as determined by the support of $L_{\theta}$) and an optimal classical estimator for post-processing data, such as the maximum-likelihood estimator in the limit of large data sets Olivares and Paris (2009); Pezzé et al. (2007); Hradil et al. (1996); Le Cam (1986); Cramer (1946); Braunstein and Caves (1994); Helstrom (1967) (refer to Refs. Rubio and Dunningham (2019, 2020) for Bayesian analysis in the scenario of limited data). The optimal POVM ensures that the CFI equals the QFI, i.e., $I(\theta)=I_{Q}(\theta)$. This is crucial as it determines that the single-parameter quantum Cramér-Rao bound can always be achieved, provided that both the estimator and the measurement basis are optimal.
The SLD operator $L_{\theta}$ is an essential ingredient in the quantum parameter estimation framework, which can readily be obtained from the solution of the Lyapunov equation as Paris (2009):
$L_{\theta}=2\int_{0}^{\infty}e^{-t\rho_{\theta}}\partial_{\theta}\rho_{\theta}e^{-t\rho_{\theta}}dt,$ (8)
@@
$L_{\theta}=2\sum_{n,m}\frac{\langle\phi_{m}|\partial_{\theta}\rho_{\theta}|\phi_{n}\rangle}{\epsilon_{n}+\epsilon_{m}}|\phi_{m}\rangle\langle\phi_{n}|$ (9)
which is obtained from Eq. (8) using the spectral decomposition of $\rho_{\theta}$, that is:
$\rho_{\theta}=\sum_{i}\epsilon_{i}|\phi_{i}\rangle\langle\phi_{i}|,$ (10)
where $\epsilon_{i}$ and $|\phi_{i}\rangle$ are the $i$-th eigenvalue and eigenvector of $\rho_{\theta}$, respectively. From Eq. (9), it is straightforward to obtain the QFI in this basis as follows:
$I_{Q}(\theta)=2\sum_{n,m}\frac{|\langle\phi_{m}|\partial_{\theta}\rho_{\theta}|\phi_{n}\rangle|^{2}}{\epsilon_{n}+\epsilon_{m}},\hskip 14.22636pt\epsilon_{n}+\epsilon_{m}\neq 0.$ (11)
@@
$I_{Q}(\lambda)=I_{\epsilon}(\lambda)+2\sum_{n\neq m}\sigma_{nm}|\langle\phi_{m}|\partial_{\lambda}\phi_{n}\rangle|^{2},$ (12)
where $I_{\varepsilon}(\lambda)$ is the (classical) Fisher information of the distribution of the eigenvalues of $\rho_{\lambda}$, and $\sigma_{nm}$ may be written as Paris (2009)
$\sigma_{nm}=\frac{(\epsilon_{n}-\epsilon_{m})^{2}}{\epsilon_{n}+\epsilon_{m}}+\text{any antisymmetric term}.$ (13)
@@
The QFI of Eqs. (6)-(11) applies to any quantum state. However, for pure quantum states $\rho_{\theta}=|\psi_{\theta}\rangle\langle\psi_{\theta}|$ such that $\text{Tr}[\rho_{\theta}^{2}]=1$, SLD $L_{\theta}=2\partial_{\theta}\rho_{\theta}$ and the QFI adopts a simpler form given by Paris (2009); Sidhu and Kok (2020); Liu et al. (2019):

### v3

@@ 2. Optimal measurement parameter-value
Quantum criticality is in general a property of equilibrium states such as the Gibbs thermal state (which becomes the ground state at zero temperature limit) Sachdev (1999); Vojta (2003). Preparing such equilibrium states can be challenging in practice, demanding ultra-low temperatures or extremely long preparation times. One can in fact analyze the ratio of the Fisher information and the time cost to discern the practical quantum advantage. The time-dependent bound on the Fisher information Boixo et al. (2007) provides valuable insights on the prediction of such advantage. As will be discussed later within a broader resource analysis segment, this leads to outcomes specific to the critical system. In addition, criticality-enhanced quantum precision is only achievable around the phase transition point. This requires fine tuning and thus is mostly beneficial in the context of local sensing, where significant prior information about the parameter of interest is available. Beyond criticalities, the sensing capability of many-body systems have also been investigated in their strongly correlated phases, even in the presence of thermal fluctuations Mihailescu et al. (2023, 2024b, 2024a). Moreover, as an alternative approach, one can also use non-equilibrium dynamics of many-body systems for sensing unknown parameters. These setups are more experimental friendly as their initial state can be a simple product state and entanglement is generated naturally during the dynamics. Non-equilibrium dynamics in many-body systems can be induced through different methods, including quantum quench Mitra (2018); Daley et al. (2012); Calabrese and Cardy (2006); Calabrese et al. (2011); Shi et al. (2024); Gietka et al. (2022), measurement quench Pouyandeh et al. (2014); Bayat et al. (2018); Skinner et al. (2019a); Li et al. (2018a), external driving Chu and Telnov (2004); Gritsev and Polkovnikov (2017); Oka and Kitamura (2019) and adiabatic evolution Mihailescu et al. (2024b). Apart from the probe size, in non-equilibrium quantum sensors, time is also a key resource which plays an important role and has to be included in the definition of standard quantum limit and quantum-enhanced sensitivity. In general, Fisher information scales as $I\sim t^{\alpha}L^{\beta}$, where $t$ is time, $L$ is the system size and $\alpha$ and $\beta$ are two exponents Ilias et al. (2022). In non-equilibrium sensors, the standard quantum limit is defined as $\alpha{=}\beta{=}1$. Any situation with $\alpha>1$ or $\beta>1$ is called quantum-enhanced sensitivity with the special case of Heisenberg limit, defined as $\alpha{=}\beta{=}2$. In certain systems in which the unitary dynamics preserves the number of excitations, the Fisher information may scale with the number of excitations too. In fact, in Ref. Manshouri et al. (2024) it was shown that in excitation-preserving dynamics Fisher information scales as $I\sim t^{\alpha}L^{\beta}N_{\rm ex}^{\gamma}$, where $N_{\rm ex}$ is the number of excitations and $\gamma$ is its corresponding exponent. This is an interesting observation which shows how the sensitivity increases from single excitation subspace, i.e. $N_{\rm ex}{=}1$, to half filling, i.e. $N_{\rm ex}{=}L/2$. In addition, Ref. Gietka and Ritsch (2023) proposes using spin-orbit coupled quantum gases which they claim can be utilized to overcome the Heisenberg limit, and Ref. Gietka et al. (2022) demonstrates that Fisher information can exhibit exponential scaling as a physical effect by quenching dynamics far beyond the critical point. Unlike criticality-based quantum sensors, the criteria for achieving quantum-enhanced precision is not yet well-characterized in non-equilibrium quantum many-body probes. For instance, we know that localized phases are not good for sensing He et al. (2023); Yousefjani et al. (2023a, 2025a); Sahoo et al. (2024) but what features in extended phases are crucial for quantum-enhanced sensitivity are not yet identified.
@@
In practice, no quantum device is perfectly isolated from its environment and thus decoherence is an inevitable part of any quantum protocol. Consequently, any quantum dynamics is actually an open system evolution. Different formalisms have been developed for addressing open system dynamics, among them are: (i) master equation Carmichael (1993); and (ii) non-Hermitian Hamiltonians Rotter (2009). In the former, the evolution of the system and its environment is described by a unitary operation. By tracing out the environment, the evolution of the system, typically within Born-Markov approximations, is thus characterized by the Lindblad (or Gorini-Kossakowski-Sudarshan-Lindblad) master equation Breuer et al. (2002). In this formalism a quantum system initially prepared in a pure state may indeed evolve to a mixed state due to the action of jump operators — a signature of information leakage (irreversible process) to the environment. In non-Hermitian formalism, the environment is monitored for tracking a trajectory in which no jump operator action takes place Plenio and Knight (1998); Daley (2014). As a result, an initially pure quantum state remains pure throughout the dynamics, despite being evolved effectively as an open quantum system. Other mechanisms that address or profit from dissipation include decoherence-free subspaces Lidar et al. (1998); Duan and Guo (1997), by coupling the quantum resources to a common environment that can be measured at least in part Braun and Martin (2011) and coherent averaging (Fraïsse and Braun, 2015) more generally leading to quantum-enhanced measurements without entanglement (Braun et al., 2018), dynamical decoupling Viola et al. (1999); Khodjasteh and Lidar (2005), and quantum reservoir engineering Poyatos et al. (1996). The latter technique involves adding an engineered Lindbladian term to the open system dynamics. This approach can be designed to steer the system into a desired pure steady-state, effectively cooling the system to that state Myatt et al. (2000); Pielawa et al. (2007, 2010); Groszkowski et al. (2022). Both master equation and non-Hermitian formalisms have been investigated for quantum sensing. Indeed, a general joint (system plus environment) quantum Fisher information upper bound has been formulated Escher et al. (2011a). Hence, one would expect that such open quantum system dynamics may also achieve quantum-enhanced sensitivity provided that the system goes through a phase transition, which is again accompanied by an eigenvalue gap closing. In the master equation formalism, the phase transition is determined for the steady state of the system at the point that the eigenvalues of the Liouvillian operator shows gap closing Minganti et al. (2018). An example of such systems with quantum-enhanced sensitivity is boundary time crystals Montenegro et al. (2023). In the non-Hermitian systems, quantum-enhanced sensitivity has been shown to emerge at the exceptional points Wiersig (2020), where two or more eigenvalues and corresponding eigenstates coalesce, as well as in the cases of different types of gap closing in the complex energy plane Sarkar et al. (2024).
@@
In this review, we explore the capacity of many-body systems for serving as quantum sensors in both equilibrium and non-equilibrium regimes. Our paper is complementary to a few previous review papers on different aspects of quantum sensing. In Ref. Paris (2009), the theoretical aspects of quantum estimation theory are comprehensively discussed. On the other hand, Ref. Degen et al. (2017) mostly focuses on experimental implementation of quantum sensors. More recent reviews have been dedicated for exploring Fisher information and its properties Meyer (2021); Liu et al. (2020) and entanglement-enhanced metrology Huang et al. (2024a). In addition, recently an essay has also been dedicated to atomic, molecular, and optical platforms for realizing quantum sensors Ye and Zoller (2024). While quantum many-body sensors have received significant attention in recent years, these aforementioned recent progresses have not been covered in the previous review articles. This makes it timely to provide a review on the progress towards quantum many-body sensors.
@@ II Elements of Parameter Estimation
Precise measurements of relevant quantities are essential in science. In physics, high-precision measurements lead us to test the theory of special relativity Einstein (1923, 1905); Michelson and Morley (1887); Rossi and Hall (1941); Hafele and Keating (1972a, b), scrutinize the theory of general relativity via gravitational wave detection Einstein (1923); Phy (2016), and validate the Standard Model of physics by confirming the existence of the elusive Higgs particle Englert and Brout (1964); Higgs (1964); Guralnik et al. (1964); The CMS Collaboration (2022). Thus, pursuing novel schemes and strategies for achieving high-precision measurements are highly desirable for advancing science. However, measuring a relevant quantity is only sometimes a straightforward procedure. Furthermore, quantities such as coupling strengths, temperature, and in general, those lacking corresponding quantum observables, such as entanglement Genoni et al. (2008); Brida et al. (2010) or purity Brunelli et al. (2012), can only be determined indirectly. Therefore, one requires to infer a relevant quantity of the system by indirect means. The art of inferring such a relevant quantity by means of measuring another is known as parameter estimation Sidhu and Kok (2020); D’ariano et al. (2000); Paris (2009)—see recent studies on the connection between quantum speed limits and quantum parameter estimation Maleki et al. (2023); Herb and Degen (2024); del Campo et al. (2013). For a comprehensive review of this relationship, see Ref. Deffner and Campbell (2017).
@@
Figure 2: Sketch of parameter estimation steps: Standard scheme for estimating an unknown parameter $\theta$. An initial quantum state $\rho$ dynamically encodes the parameter $\theta$ to be estimated as $\rho(\theta)$. To extract information about $\theta$, a set of measurements via positive operator-valued measure (POVM) $\{\hat{\Pi}_{x}\}$, are performed on the probe with random outcomes $x$. The resulting measurement outcomes are fed into a classical estimator $\tilde{\Theta}$ to infer the unknown parameter $\theta$.
All the above steps need optimization to successfully estimate the parameters. It is important to note that the actual values of the parameters are always unknown by definition, and one must infer them indirectly. One common methodology to quantify the uncertainty in estimating a single unknown parameter $\theta$ is accessing the system of interest through measurements with observable $\hat{O}$, which leads to estimating the error of $\theta$ using the error propagation formula Sidhu and Kok (2020):
$\delta\theta=\frac{\Delta\hat{O}}{\left|\partial_{\theta}\langle\hat{O}\rangle\right|},\hskip 28.45274pt\frac{\partial}{\partial\theta}:=\partial_{\theta}$ (1)
where $\Delta\hat{O}$ and $\langle\hat{O}\rangle$ are the standard deviation and the expectation value of the outcomes associated with $\hat{O}$, respectively. The denominator in Eq. (1) indicates that in order to decrease the error in the estimation of $\theta$ (i.e., $\delta\theta\ll 1$), one needs to increase the denominator, i.e., $\left|\partial_{\theta}\langle\hat{O}\rangle\right|\gg 1$. To achieve this, it is necessary to find a set of outcomes with an expectation value $\langle\hat{O}\rangle$ that varies significantly with respect to $\theta$. Conversely, a set of outcomes that varies weakly with $\theta$ (i.e., flat slope $|\partial_{\theta}\langle\hat{O}\rangle|\ll 1$) will result in a substantial error (i.e., $\delta\theta\gg 1$). While Eq. (1) offers a heuristic explanation of a necessary ingredient for estimating a parameter, a more rigorous and operational framework must be provided to determine a formal figure of merit for quantifying estimation performance. In fact, Eq. (1) does not take into account the entire probability distribution of the outcomes and cannot provide the framework to determine a proper figure of merit for quantifying the performance of an estimation strategy.
@@ II.1 Global and local estimation theory
A suitable strategy to estimate a parameter usually depends on the amount of a priori information one has on the parameter itself. If no a priori information is available, the estimation strategy, i.e. the choice of the measure to be performed on the probe and the procedure employed to process data, should be optimal in average i.e. should be equally good for any value of the parameter. The set of techniques and methods used to choose and optimize the figure of merit in this case is known as global estimation theory Helstrom (1976); Holevo (2011a), see Ref. Mukhopadhyay et al. (2025) for a recent review on global estimation theory. Once a global estimation has been performed an experimenter is left with some a priori information to be improved, i.e. one should look for figures of merits that select the best estimation strategy for a specific value or range of values of a parameter. This optimization problem is addressed by local estimation theory Bengtsson and Życzkowski (2017); Van Trees (2004); Helstrom (1976) which allows one to achieve the ultimate bound on precision. The quantum version, namely the local quantum estimation theory Braunstein and Caves (1994); Paris (2009) is the fundamental tool of quantum probing, sensing and metrology and is the main subject of the present review, with the exception of Section XIV, where various aspects of global estimation theory are discussed.
@@ II.3 Optimizing the classical bound: the QFI
The conditional probability of obtaining the outcome $x$ for a given $\theta$ obeys the Born rule:
$p(x|\theta)=\text{Tr}[\hat{\Pi}_{x}\rho(\theta)].$ (4)
Assuming regularity conditions, and since the trace and the derivative are linear operators, the derivative of the probability distribution $\partial_{\theta}p(x|\theta)$ translates into taking the derivative of the quantum state $\partial_{\theta}\rho(\theta)$, specifically $\partial_{\theta}p(x|\theta)=\text{Tr}[\hat{\Pi}_{x}\partial_{\theta}\rho(\theta)]$. Here, the derivative of the quantum state satisfies the Lyapunov equation given by:
$\partial_{\theta}\rho(\theta)=\frac{1}{2}\{\hat{L}(\theta),\rho(\theta)\},$ (5)
@@
$I(\theta)\leq I_{Q}(\theta):=\text{Tr}[\rho(\theta)\hat{L}(\theta)^{2}].$ (6)
$I_{Q}(\theta)$ denotes the quantum Fisher information (QFI), leading to the quantum Cramér-Rao bound Paris (2009); Yuen and Lax (1973); Helstrom and Kennedy (1974); Braunstein and Caves (1994); Sidhu and Kok (2020); Albarelli et al. (2020a); Liu et al. (2019):
@@
Note that the derivative of the quantum state satisfies other expressions, e.g., the Right Logarithmic Derivative (RLD) Albarelli et al. (2020a). This gives rise to different quantum lower bounds— see Ref.Meyer (2021) for a thorough discussion on the uniqueness of CFI and QFI, and Ref.Albarelli et al. (2020a) for an overview of different achievable bounds.
The QFI of Eq. (6) accounts for the ultimate precision limit for the estimation of $\theta$. Note that the QFI, as the maximization of the CFI over all possible measurements, depends on the quantum statistical model $\rho(\theta)$ and the SLD $\hat{L}(\theta)$ only, and it does not depend explicitly on a particular measurement basis. The optimal POVM can be composed with the set of projectors over the eigenstates of $\hat{L}(\theta)$ with an associated optimal quantum estimatorParis (2009).
To saturate the quantum Cramér-Rao bound in Eq. (2), one needs both the optimal POVM (as determined by the support of $\hat{L}(\theta)$) and an optimal classical estimator for post-processing data, such as the maximum-likelihood estimator in the limit of large data sets Olivares and Paris (2009); Pezzé et al. (2007); Hradil et al. (1996); Le Cam (1986); Cramer (1946); Braunstein and Caves (1994); Helstrom (1967) (refer to Refs. Rubio and Dunningham (2019, 2020) for Bayesian analysis in the scenario of limited data). The optimal POVM ensures that the CFI equals the QFI, i.e., $I(\theta)=I_{Q}(\theta)$. Thus, in practical scenarios, any observable measurement used to extract information about an unknown parameter must be compared against the fundamental sensing benchmark set by the QFI (For example, see Refs. Troiani and Paris (2018); Montenegro (2025) for a theoretical demonstration of QFI saturation via magnetization measurements and near-saturation through homodyne techniques Montenegro et al. (2020), as well as Ref. Yu et al. (2022) for experimental work achieving near QFI saturation in the phase estimation of a solid-state spin system). Crucially, the optimal POVM always exists and plays a key role as it guarantees that the quantum Cramér-Rao for single-parameter estimation can always be attained, provided that both the estimator and the measurement basis are are chosen optimally.
The SLD operator $\hat{L}(\theta)$ is an essential ingredient in the quantum parameter estimation framework, which can readily be obtained from the solution of the Lyapunov equation as Paris (2009):
$\hat{L}(\theta)=2\int_{0}^{\infty}e^{-t\rho(\theta)}\partial_{\theta}\rho(\theta)e^{-t\rho(\theta)}dt,$ (8)
@@
$\hat{L}(\theta)=2\sum_{n,m}\frac{\langle\phi_{m}|\partial_{\theta}\rho(\theta)|\phi_{n}\rangle}{\epsilon_{n}+\epsilon_{m}}|\phi_{m}\rangle\langle\phi_{n}|$ (9)
which is obtained from Eq. (8) using the spectral decomposition of $\rho(\theta)$, that is:
$\rho(\theta)=\sum_{i}\epsilon_{i}|\phi_{i}\rangle\langle\phi_{i}|,$ (10)
where $\epsilon_{i}$ and $|\phi_{i}\rangle$ are the $i$-th eigenvalue and eigenvector of $\rho(\theta)$, respectively. From Eq. (9), it is straightforward to obtain the QFI in this basis as follows:
$I_{Q}(\theta)=2\sum_{n,m}\frac{|\langle\phi_{m}|\partial_{\theta}\rho(\theta)|\phi_{n}\rangle|^{2}}{\epsilon_{n}+\epsilon_{m}},\hskip 14.22636pt\epsilon_{n}+\epsilon_{m}\neq 0.$ (11)
@@
$I_{Q}(\theta)=\bar{I}(\theta)+2\sum_{n\neq m}\epsilon_{nm}|\langle\phi_{m}|\partial_{\theta}\phi_{n}\rangle|^{2},$ (12)
where $\bar{I}(\theta)=\sum_{i}(\partial_{\theta}\epsilon_{i})^{2}/\epsilon_{i}$ is the CFI of the distribution of the eigenvalues of $\rho(\theta)$, and $\epsilon_{nm}$ may be written as Paris (2009)
$\epsilon_{nm}=\frac{(\epsilon_{n}-\epsilon_{m})^{2}}{\epsilon_{n}+\epsilon_{m}}+\text{any antisymmetric term}.$ (13)
@@
The QFI of Eqs. (6)-(11) applies to any quantum state. However, for pure quantum states $\rho(\theta)=|\psi(\theta)\rangle\langle\psi(\theta)|$ such that $\text{Tr}[\rho(\theta)^{2}]=1$, SLD $\hat{L}_{\theta}=2\partial_{\theta}\rho(\theta)$ and the QFI adopts a simpler form given by Paris (2009); Sidhu and Kok (2020); Liu et al. (2019):